Circular No. 8813

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2007V, 2007W, 2007X, 2007Y, 2007Z Five apparent supernovae have been discovered on unfiltered CCD

frames and reported to the Central Bureau: 2007V by N. Joubert and W. Li (LOSS/KAIT; cf. IAUC 8804); 2007Z by E. J. Christensen (Mt. Lemmon Survey 1.5-m telescope); and the rest by L. A. G. Monard (Pretoria, S. Africa, 0.35-m reflector). Discovery observations:

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007V Feb. 12.46 10^{h}27^{m}05.^{s}58 − 3^{o}18′05′′.1 17.9 0′′.3 E, 1′′.8 S
2007W Feb. 15.037 13 21 48.50 −13 12 56.9 17.1 8′′ W, 33′′ S
2007X Feb. 15.068 14 30 14.50 −36 57 53.1 15.3 30′′ E, 2′′ S
2007Y Feb. 15.774 3 02 35.92 −22 53 50.1 17.5 24′′ W, 110′′ S
2007Z Feb. 17.221 5 15 42.68 +19 21 49.5 17.5 10′′.4 W, 2′′.7 N

Additional information is available on these objects on the following Electronic Telegrams: 2007V, CBET 841 (type Ia, discovered ∼ 1 month past maximum); 2007W in NGC 5105, CBETs 844 and 846 (type II); 2007X in ESO 385-G32, CBETs 844 and 849 (young type II); 2007Y in NGC 1187, CBET 845; 2007Z in PGC 16993, CBETs 847 and 849 (young type II). Additional LOSS/KAIT magnitudes for 2007Y from R. Mostardi and W. Li: 2006 Dec. 24.27 UT, [18.1; 2007 Jan. 19.23, [17.5; 26.21, [17.3; Feb. 15.15, 16.7 (independent discovery). Li provides position end figures 35s.94, 49”.4 for 2007Y.

COMETS C/2006 W5, C/2006 W6, C/2006 W7, C/2006 X2 (SOHO)

Additional Kreutz sungrazing comets have been found on SOHO website images (cf. IAUC 8811). K. Battams notes that C/2006 W5 appeared as a fuzzy streak ∼ 2′ long and quite faint (mag ∼ 7.5). C/2006 W6 was very small, quite diffuse, and very faint (mag 8). C/2006 W7 was also very small and faint (mag ∼ 7.5), and slightly diffuse. C/2006 X2 was very diffuse, faint, and somewhat elongated.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 W5 Nov. 20.004 15^{h}40.^{m}3 −21^{o}16′ C2 TH 2007-B76
C/2006 W6  20.038 15 40.2 −21 21 C2 TH 2007-B76
C/2006 W7  25.479 16 03.1 −22 28 C2 RK 2007-B76
C/2006 X2 Dec. 9.049 17 04.6 −24 41 C2 RK 2007-B76
2007 February 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green